### v1

# MILD: Multi-Layer Diffusion Strategy for Complex and Precise Multi-IP Aware Human Erasing

Abstract
Recent years have witnessed the remarkable success of diffusion models, especially in various image-customized tasks. Previous works have achieved notable advances on simple human-oriented erasing task by leveraging explicit mask guidance and semantic-aware inpainting paradigm. Despite this progress, existing methods still struggle with significant human removal challenges raised by “Multi-IP Interactions” under more complex realistic scenarios such as human-human occlusions, human-object entanglements and human-background interferences. These approaches are typically limited by: 1) Dataset deficiency. They basically lack large-scale multi-IP datasets, especially those covering complex scenarios such as dense occlusions, camouflaged or distracted backgrounds and diverse IP interactions. 2) Lack of spatial decoupling. Most methods lack effective spatial decoupling strategies to disentangle different foreground instances from the background, which limits their ability to achieve targeted human removal and clear background inpainting. These limitations result in performance degradation in precisely erasing overlapping individuals and restoring occluded regions in real-world scenes. In this work, we first introduce a high-quality human erasing dataset, MILD dataset, capturing diverse pose variations, occlusions, and complex backgrounds in multi-IP interactive scenarios. Building on this foundation, we propose Multi-Layer Diffusion (MILD), a novel multi-layer diffusion strategy for complex and precise multi-IP aware human erasing. Specifically, MILD first decomposes the traditional generation process into semantically separated denoising pathways, enabling independent reconstruction of each foreground instance and the background. To enhance human-centric understanding, we introduce Human Morphology Guidance, a plug-and-play guidance module that integrates pose, parsing, and spatial relations into the generation process, improving the morphological awareness and promoting more effective restoration. We further present Spatially-Modulated Attention, an adaptive mechanism that leverages spatial mask priors to effectively modulate attention across semantic regions, resulting in fewer boundary artifacts and mitigated semantic leakage. Extensive experiments demonstrate that MILD consistently outperforms state-of-the-art methods on challenging human erasing benchmarks. Our project page is available at: https://mild-multi-layer-diffusion.github.io/.
{strip}
Figure 1: The figure exhibits MILD (our method) can robustly handle three critical challenges in human erasing: Human–Human Occlusion, Human–Object Entanglement, and Human–Background interferences and can achieve clean and artifact-free removal results across diverse scenarios. Note each case shows the input and corresponding output pair.
1 Introduction
Object removal aims to eliminate user-specified regions and synthesize visually coherent, context-consistent backgrounds Jampani et al. (2021); Xu et al. (2023). This capability supports a wide range of applications, including photo editing, privacy preservation, and interactive content creation Yu et al. (2018); Liu et al. (2018). However, achieving both thorough removal and faithful background restoration remains a challenging task, especially in scenarios involving multiple interacting instances and dense occlusions.
Early non-parametric methods Criminisi et al. (2004); Barnes et al. (2009) repainted erased regions by copying patches from visible areas, but often produce repetitive textures and fail in complex scenes. Afterwards, GAN-based models Pathak et al. (2016); Iizuka et al. (2017); Yu et al. (2018); Nazeri et al. (2019) introduce adversarial learning to enhance visual realism, yet suffer from global inconsistency and instability under large masks. Recently, diffusion-based models Song et al. (2021); Avrahami et al. (2023); Yang et al. (2023a) have achieved competitive performance by leveraging large-scale pretrained generative priors Rombach et al. (2022b); Ekin et al. (2024). These models Zhuang et al. (2024); Yildirim et al. (2023) excel in generating diverse content, but remain limited in eliminating objects in complex multi-instance scenarios. For example, in Figure 4, standard LDM always fails to remove the target cleanly and instead hallucinates a new human-like figure at the masked region. Notably, the generated content retains dominant visual hints, such as “human”, implying that semantic features from the removed subject have leaked into the generation process.
These limitations highlight a fundamental challenge: most existing methods treat object removal as an unified generation task, without explicitly decoupling the generation pathways for different foreground instances. This framework often leads to semantic interference and content entanglement, especially in multi-IP or object-dense scenes, where features from one instance may bleed into others and degrade both the quality and controllability of the output.
To address these challenges, we introduce Multi-Layer Diffusion (MILD), a diffusion-based strategy that reformulates human erasing as a disentangled, layer-wise, instance-aware generation process. As illustrated in Figure 1, MILD achieves high-fidelity removal in three representative challenges: human-human occlusions, human-object entanglements, and human-background interferences, which represent typical scenarios where existing methods tend to perform suboptimally. Our contributions are threefold:
  • MILD dataset. A high-quality human erasing dataset is curated capturing diverse scenes with challenging poses, occlusions, and human-instance interactions. This dataset provides paired images with precise removal masks and realistic background, enabling rigorous evaluation of human erasing models in real-world scenarios.
  • Multi-Layer Diffusion Strategy (MILD). We propose a MILD backbone that decomposes generation process into multiple semantically separated LoRA-based generation pathways, producing composable outputs corresponding to masked instances and the background. To further enhance performance, we introduce Human Morphology Guidance for instance-level understanding and Spatially-Modulated Attention for spatially conditioned attention control, enabling precise removal with minimal semantic leakage in complex scenes.
  • Comprehensive Experiments. Extensive experiments on both real-world and synthetic datasets evaluate the effectiveness of our MILD strategy using diverse quantitative metrics and human evaluations. The ablation studies further demonstrate the individual contributions of Human Morphology Guidance and Spatially-Modulated Attention. Collectively, the results confirm that MILD achieves state-of-the-art performance in removal accuracy and background fidelity, outperforming existing human erasing baselines.
Figure 2: (a) Overview of MILD. Given an input image and a set of target masks, the proposed Multi-Layer Diffusion (MILD) strategy performs human erasing by generating disentangled foreground layers and a background layer. A shared UNet with Layered LoRA enables efficient denoising across all branches, producing composable outputs. Spatially-Modulated Attention (SMA) injects adaptive biases to suppress semantic leakage across masked regions. Cross-attention is conditioned on the text prompt and Human Morphology Guidance (HMG) to guide instance-aware generation. (b) HMG extracts fine-grained human priors from pose and parsing maps using two dedicated encoders. These features, combined with a spatial mask prior, form a unified latent representation that enriches the denoising process with human-centric and spatial cues, improving fidelity and identity consistency.
2 Related Work
Image Inpainting
Image inpainting aims to restore missing or occluded regions in a visually coherent and semantically consistent manner in various scenarios Sargsyan et al. (2023); Zhou et al. (2020); Wang et al. (2024); Yoon and Cho (2024). Traditional methods relied on patch propagation Hays and Efros (2007); Ding et al. (2019), transferring low-level textures from visible regions but failing in semantically complex scenes. With the rise of deep learning, the encoder-decoder Zhou et al. (2023); Cao et al. (2023); Altinel et al. (2018); Li et al. (2020a); Yi et al. (2020a) and GAN-based architectures Iizuka et al. (2017); Yu et al. (2018); Zeng et al. (2019) have significantly improved structural plausibility and perceptual quality, further enhanced by edge guidance Nazeri et al. (2019) and attention mechanisms Yu et al. (2018). Diffusion-based methods Lugmayr et al. (2022); Saharia et al. (2022) have recently achieved strong generation performance by operating in latent spaces guided by powerful pretrained priors, but they often lack precise spatial control or instance-level disentanglement. Methods like SDXL Inpainting Podell et al. (2023) introduce mask-based conditioning for edit localization, but treating all masked regions as single entity prevents instance-specific control. Meanwhile, recent efforts incorporate large language models (LLMs) to enable instruction-based editing Brooks et al. (2023); Huang et al. (2024); Li et al. (2024b); Sheynin et al. (2024); Li et al. (2024c), yet they struggle to localize individual instances in complex scenes.
Object Removal
Object removal, as a task derived from image inpainting, aims to eliminate user-specified content while preserving scene fidelity. Early approaches Suvorov et al. (2022) achieved reasonable results by adapting general-purpose generative models with task-specific Jain et al. (2023); Yi et al. (2020b) guidance, but they struggled to capture complex textures and scene structures. Recent works Hu et al. (2022); Kingma and Welling (2014); Ronneberger et al. (2015); Rombach et al. (2022a); Ho et al. (2020) have explored a range of strategies to improve controllability and fidelity Winter et al. (2024). Training-based Yang et al. (2023b); Jiang et al. (2025); Liu et al. (2025); Li et al. (2025) approaches fine-tune diffusion models using supervision such as masks, prompts, or parameter-efficient modules. Training-free Sun et al. (2025); Han et al. (2024); Jia et al. (2024); Li et al. (2024a) approaches, in contrast, manipulate internal representations (e.g., attention maps or embeddings) during inference to enable flexible editing. In parallel, different forms of user interaction have been introduced—ranging from text-prompted editing Bar-Tal et al. (2022); Zhao et al. (2021); Geng et al. (2024) to mask-guided Li et al. (2022); Köhler et al. (2014); Xie et al. (2023a, b) generation—to support more intuitive control. Despite these advances, we found that most existing methods lack instance-level distinction, generating a single prediction conditioned on all masked targets or prompts. This entangled treatment limits the compositional flexibility and often leads to interference between each instance and its surrounding region.
3 Method
Overview
Existing inpainting methods often fail to entirely remove a instance, leaving residual artifacts or introducing semantic leakage in complex scenes. To address these limitations, we introduce Multi-Layer Diffusion (MILD) (as shown in Figure 2), a diffusion-based strategy that reformulates human erasing as a layered generation process. MILD features a structured extension of low-rank adaptation as its backbone (Section 3.1) to produce disentangled per-instance and background layers. On top of this, Human Morphology Guidance (Section 3.2) injects pose, parsing, and spatial priors to enhance human-aware understanding, while Spatially-Modulated Attention (Section 3.3) dynamically suppresses semantic leakage across spatially separated regions with mask priors. Together, these components enable precise, artifact-free erasing in scenes with occlusions and entanglements.
3.1 MILD Backbone: Layered LoRA Diffusion
Traditional inpainting methods usually struggle with multi-IP scenes, as their unified generation process tends to produce residual artifacts and semantic confusion due to the lack of instance-level disentanglement. To address this limitation, we introduce a novel MILD backbone, a multi-branch adaptation of Low-Rank Adaptation (LoRA) that enables disentangled reconstruction of each instance while ensuring coherent scene composition.
Given an input image and $N$ target masks $\{M_{k}\}_{k=1}^{N}$, MILD Backbone produces $N$ foreground layers and a clean background layer. Each foreground branch is conditioned on its corresponding mask and reconstructs the associated subject, while the background branch synthesizes the unobstructed environment with all instances removed. To avoid redundant parameters and enable scalability, we adopt a shared UNet backbone with LoRA adapters. A single set of low-rank parameters is used across all foreground branches, while the background branch maintains an independent set. Formally, each LoRA modification is defined as:
$\Delta W_{p}^{(b)}$ $=\alpha_{p}^{(b)}\,B_{p}^{(b)}\,A_{p}^{(b)},$ $p\in\{Q,K,V\},b\in\{\mathrm{fg},\mathrm{bg}\},$ (1)
where $A_{p}^{(b)}\in\mathbb{R}^{r\times d}$, $B_{p}^{(b)}\in\mathbb{R}^{d\times r}$, and $\alpha_{p}^{(b)}$ is a learnable scaling factor that supports concurrent training and inference of all branches, and facilitates interaction between instance-level predictions and global background restoration.
During training, each foreground branch learns to reconstruct its assigned region by utilizing the loss:
$\mathcal{L}_{k}=\left\|M_{k}\odot\left(\epsilon-\epsilon_{k}\right)\right\|_{2}^{2},$ (2)
where $\epsilon\sim\mathcal{N}(0,1)$ is standard Gaussian noise and $\epsilon_{k}$ denotes the prediction from the $k$-th branch. The background branch predicts noise over the entire image using the loss:
$\mathcal{L}_{\text{bg}}=\left\|\epsilon-\epsilon_{\text{bg}}\right\|_{2}^{2}.$ (3)
The total loss is formulated by combining the foreground and background losses as follows:
$\mathcal{L}_{\text{total}}$ $=\lambda_{t}\sum_{k=1}^{N}\mathcal{L}_{k}+\mathcal{L}_{\text{bg}},$ (4)
[…]
where $\lambda_{t}$ gradually increases to stabilize training. We adopt a staged strategy, first optimizing the background branch before introducing foreground supervision through ramp-up.
During inference, MILD produces composable outputs:
$\bigl{\{}\text{Layer}_{1},\ldots,\text{Layer}_{N},\text{Background}\bigr{\}},$ (5)
where each $\text{Layer}_{k}$ represents an individual instance $x_{k}$. These layers can be selectively combined to reconstruct arbitrary subsets by removing the instance in $\mathcal{R}\subseteq\{1,2,\ldots,N\}$ as:
$x_{\text{final}}=\Bigl{(}\sum_{k\notin\mathcal{R}}M_{k}\odot x_{k}\Bigr{)}+\Bigl{(}\Bigl{(}1-\sum_{k\notin\mathcal{R}}M_{k}\Bigr{)}\odot x_{\text{bg}}\Bigr{)},$ (6)
where $\mathcal{R}$ denotes the set of instances to remove. For each $k\in\mathcal{R}$, the corresponding layer is excluded from the composite and automatically filled by the background layer $x_{\text{bg}}$. This layered formulation supports controllable recomposition and improves reconstruction fidelity through joint foreground-background supervision.
Algorithm 1 MILD Training
0: Image input: $x$, masks $\{M_{k}\}_{k=1}^{N}$
0: Condition inputs: text prompt $c_{\text{text}}$, pose keypoints $P$, parsing maps $S$
1:// Human Morphology Guidance
[…]
4:for $k=1$ to NN do
5:  $M_{\text{others}}^{(k)}\leftarrow\text{Aggregate}(\{M_{j}\}_{j\neq k})$
6:  $Z_{\text{mask}}^{(k)}\leftarrow\text{MaskEncoder}(M_{\text{others}}^{(k)})$
7:  $C^{(k)}\leftarrow\text{Concat}(Z_{\text{text}},Z_{\text{pose}},Z_{\text{parse}},Z_{\text{mask}}^{(k)})$
8:end for
[…]
10:// Dual LoRA Training
11:for each timestep $t$ and noise $\epsilon$ do
12:  $z_{t}\leftarrow\sqrt{\bar{\alpha}_{t}}z+\sqrt{1-\bar{\alpha}_{t}}\epsilon$
13:  for $k=1$ to NN do
14:   $\mathcal{L}_{k}\leftarrow\|(1-M_{k})\odot(\epsilon-\epsilon_{k})\|_{2}^{2}$ {Foreground branch}
15:  end for
16:  $\mathcal{L}_{\text{bg}}\leftarrow\|\epsilon-\epsilon_{\text{bg}}\|_{2}^{2}$ {Background branch}
17:  $\mathcal{L}_{\text{total}}\leftarrow\lambda_{t}\sum_{k=1}^{N}\mathcal{L}_{k}+\mathcal{L}_{\text{bg}}$
18:end for
3.2 Human Morphology Guidance
Although the dedicated Layered LoRA diffusion backbone lays the foundation for disentangling instance layers, fine-grained instance awareness remains essential for accurately removing targets and preserving body structures in spatially overlapping scenarios. To handle such complex scenarios, we introduce Human Morphology Guidance (HMG), a plug-and-play module that introduces fine-grained, human-centric priors into the denoising process.
As illustrated in Algorithm 1, HMG is seamlessly integrated into the MILD backbone by embedding pose and parsing information into the conditioning signals for both foreground and background branches. Specifically, preprocessed posure and parsing maps are separately encoded, flattened, linearly projected, and concatenated to form comprehensive human features, as shown in Figure 2 (b).
To further enhance spatial reasoning, we introduce a lightweight mask-based conditioning scheme that explicitly models neighboring instances. For each instance $k$, a context mask is constructed as:
$M^{(k)}=\text{clamp}\Bigl{(}\sum_{j\neq k}M_{j},\,0,1\Bigr{)},$ (7)
which captures the spatial layout of surrounding instances, providing instance-aware guidance for precise IP localization. This mask is encoded through a shallow convolutional layer and concatenated with the target’s mask embedding and human features. Together, these spatially-aware priors provide stronger guidance for accurately separating instances, particularly in occluded scenarios.
By integrating explicit morphological and spatial cues, HMG significantly enhances instance-level disambiguation, especially in challenging scenes involving occlusions, uncommon poses, or partially visible subjects.
Figure 3: Illustration of the proposed Spatially-Modulated Attention (SMA) mechanism. For each query-key pair $(i,j)$, an spatial bias $\alpha_{st}$ is applied to the vanilla attention score $A_{ij}$, where $s=m_{i}$ and $t=m_{j}$ indicate their foreground/background status.
3.3 Spatially-Modulated Attention
The original attention mechanisms utilize the entire context to generate reasonable background, but this leads to semantic leakage due to the unrestricted interactions between foreground and background tokens. To mitigate this, we propose Spatially-Modulated Attention (SMA), a lightweight module that enforces spatially constrained attention based on mask boundaries, as shown in Figure 3.
In standard self-attention, logits are computed as:
$A_{ij}=\frac{Q_{i}K_{j}^{T}}{\sqrt{d_{k}}}.$ (8)
To control attention flow across semantic separated regions, SMA introduces a learnable bias based on the binary masks $m_{i}$ and $m_{j}$ of tokens $i$ and $j$:
$\tilde{A}_{ij}=A_{ij}+\sum_{s,t\in\{0,1\}}\alpha_{st}\,\mathds{1}[m_{i}=s,\,m_{j}=t],$ (9)
where $\alpha_{st}$ are trainable parameters that modulate attention flow between foreground and background regions. The indicator function $\mathds{1}[m_{i}=s,\,m_{j}=t]$ evaluates to $1$ when token $i$ belongs to region $s$ and token $j$ to region $t$, and 0 otherwise. For instance, $\alpha_{10}$ controls how effective a foreground token attends to the background.
Initialized with zero, these biases preserve the vanilla attention at the start and enabling adaptive spatial modulation during training. This enables explicit control over cross-region interactions without altering the attention mechanism itself. With this selective suppression, the attention flow learns to “attend to what it should” and mitigates leakage across masked boundaries. Integrated with MILD backbone and HMG, SMA improves the semantic isolation and artifacts removal in challenging multi-IP scenes.
Figure 4: Qualitative results produced by MILD (ours) and other methods in real-world scenes. The masked regions (in red) and the corresponding removal results (in green) are highlighted for comparison.
4 Experiments
Method MILD Dataset OpenImages Dataset
[…]
SDXL Inpaint 53.45 0.242 0.8398 0.8331 18.22 26.22 0.1346 0.8741 0.8751 21.88
PowerPaint 48.04 0.207 0.8901 0.8668 20.27 10.56 0.0488 0.9649 0.9678 31.04
Inst-Inpaint 11.42 0.4100 0.7400 0.8207 23.75
LaMa 35.02 0.164 0.8909 0.8872 24.13 10.38 0.0470 0.9192 0.9293 31.83
CLIPAway 48.86 0.230 0.8647 0.8605 19.49 22.73 0.1283 0.8973 0.8985 23.59
RoRem 40.41 0.196 0.8888 0.8810 22.35 18.23 0.0982 0.9095 0.9148 26.59
MILD (Ours) 24.20 0.093 0.9703 0.9499 26.24 17.86 0.0668 0.9829 0.9700 32.14
Table 1: Quantitative comparison of MILD and other methods on MILD and OpenImages datasets. Best results are bold, second best are underlined. Inst-Inpaint is excluded from the evaluation on MILD dataset as it lacks mask guidance, making it unsuitable for precise localization in multi-IP scenarios.
4.1 Experimental Setup
Datasets
We conduct comprehensive evaluations on two datasets to assess both task-specific performance and cross-domain generalization: 1) Our MILD Dataset: We construct a high-quality dataset specifically tailored for multi-IP erasure. The dataset comprises $10,000$ image pairs featuring complex human interactions, diverse body poses, challenging occlusions, and a wide range of background contexts. The detailed information is introduced in Appendix A. 2) OpenImages Dataset: To evaluate generalization beyond human-centric scenarios, we randomly sample 1,000 images from OpenImages V5 Kuznetsova et al. (2020), covering diverse object categories and photographic conditions.
Implementation Details
We train MILD LoRA on 20,00020{,}000 images, including 10,00010{,}000 real-world pairs from our curated MILD dataset and 10,00010{,}000 synthetic samples generated by compositing segmented humans onto OpenImages backgrounds with random cropping and mask augmentation. The baseline model Podell et al. (2023) is optimized using AdamW Loshchilov and Hutter (2019) with a learning rate of 1e1e66, batch size $4$, and trained for 20,00020{,}000 iterations. Human Morphology Guidance leverages pose maps from OpenPose Cao et al. (2017) and parsing maps from SCHP Li et al. (2020b), preprocessed prior to training. A staged training scheme is adopted: the background branch is trained first, followed by a linear ramp-up of foreground losses after 8,0008{,}000 steps. The model employs a UNet backbone within a latent diffusion framework, equipped with rank-$16$ LoRA modules (scaling factor $16$). Inference uses $25$ DDIM steps at $512\times 512$ resolution. All experiments are conducted on NVIDIA $A800$ GPUs.
Baseline Comparisons
We compare against a diverse set of state-of-the-art object removal methods, including SDXL Inpainting Podell et al. (2023), LaMa Suvorov et al. (2022), PowerPaint Zhuang et al. (2024), Inst-Inpaint Yildirim et al. (2023), CLIPAway Ekin et al. (2024), and RoRem Li et al. (2025), covering paradigms from mask-guided inpainting to instruction-based and vision-language-guided editing. To ensure a fair comparison, we adopt the official open-source implementations and reproduce results under a unified evaluation pipeline. All models are either used as publicly released or retrained with default settings on our datasets to ensure fair comparison.
Evaluation Metrics
We adopt five complementary metrics to evaluate object removal quality from perceptual, semantic, and low-level fidelity perspectives. FID Heusel et al. (2017) evaluates global visual realism by comparing feature distributions of generated and real images, while LPIPS Zhang et al. (2018) focuses on local perceptual similarity based on deep features. DINO cosine similarity Caron et al. (2021) and CLIP Image Similarity Radford et al. (2021) measure semantic consistency by computing cosine similarity between the embeddings of generated and ground-truth images, capturing structural and high-level semantic alignment, respectively. Finally, PSNR Hore and Ziou (2010) quantifies pixel-level fidelity within the removed regions.
4.2 Quantitative and Qualitative Comparisons
Table 1 summarizes the quantitative performance of various object removal methods on the MILD and OpenImages datasets. Our proposed MILD strategy consistently outperforms all baselines on the more challenging MILD dataset and remains highly competitive on OpenImages. Compared to the SDXL Inpainting baseline, MILD reduces FID and LPIPS by $54.53\%$ and $61.57\%$ respectively, reflecting significant improvement in visual realism and perceptual quality. It also achieves higher DINO, CLIP, and PSNR scores, demonstrating enhanced semantic consistency and pixel-level fidelity through its spatially-aware layered diffusion design. While LaMa performs well on FID and LPIPS in relatively simple scenes due to its smooth outputs, it underperforms in semantic metrics like DINO and CLIP, revealing limitations in understanding complex content. In contrast, MILD delivers superior results across all dimensions, particularly in complex human interactive scenes.
Figure 4 exhibits qualitative comparisons in challenging multi-IP scenarios. When the masked human occupies a large image region (Row 1), MILD generates semantically coherent and visually realistic content outperforming baselines. In scenes with complex human interactions (Row 3), it produces cleaner boundaries and fewer redrawings. Compared to existing methods, MILD effectively reduces hallucinations, avoids oversmoothing and blurring, and reconstructs sharper details in occluded regions. Notably, while LaMa performs well in simpler cases, it struggles with complex semantics and often introduces unrealistic blending (Row 2), which consistent with our quantitative results. These results demonstrate the effectiveness of MILD’s layered diffusion strategy and spatially guided conditioning.
While traditional quantitative metrics provide valuable evaluations of model performance, they often fail to capture the direct naturalness and visual quality as perceived by human observers. To address this limitation, we conduct a comprehensive perceptual evaluation from both AI and human perspectives. The evaluation of both assessments are summarized in Figure 6, and the detailed experiment setting and the full table is shown in Appendix. Both AI and human evaluations consistently demonstrate that MILD achieves superior perceptual quality compared to baseline methods. Our method (MILD) achieves the highest success rate for both AI $(82\%)$ and human raters $(78\%)$, indicating consistent perceptual superiority. This comprehensive perceptual assessment validates the effectiveness of our approach in producing human-preferred removal results.
In summary, MILD achieves state-of-the-art performance on complex human interaction scenes, while also demonstrating competitive generalization to broader domains. It effectively balances visual realism, semantic alignment and perceptual quality across diverse human erasing scenarios.
Figure 5: Qualitative ablation comparison of our method.
Figure 6: Comparison of AI and human success rates for object removal across different methods on the MILD dataset.
4.3 Ablation Study
To evaluate the impact of the components in our multi-layer diffusion strategy, we conduct a step-by-step ablation study with both the quantitative and qualitative results in Table 2 and Figure 5. We progressively modify the baseline through four main experiments. We first incorporate the MILD Backbone into the SDXL inpainting baseline. Then the HMG and SMA modules are separately added into the structure. Finally, we evaluate the full MILD model, which integrates the MILD Backbone along with both HMG and SMA modules.
The SDXL Inpainting baseline performs poorly on the MILD dataset, often produces incomplete removals or hallucinated content when handling large or occluded subjects. Introducing MILD Backbone significantly improves both perceptual and semantic quality by disentangling generation into background and instance-specific foreground branches. This reduces the repainting phenomenon with 45.7145.71% improvement in FID metric. Building on this backbone, HMG and SMA provide complementary enhancements. As shown in Figure 5, HMG helps to preserve body structure (row $2$) and relieve the residual artifacts along object boundaries(row $1$). Meanwhile, SMA focuses on refining the attention flow, effectively suppressing the unnatural mask bleeding and background leakage, resulting in smoother transitions and more coherent backgrounds. The full MILD model combines these advantages, achieving the best overall performance across all metrics. These results validate the non-redundant and effectiveness of our modular design.
[…]
SDXL Inpaint 53.45 0.242 0.8398 0.8331 18.22
+ MILD Backbone 29.02 0.142 0.9110 0.9025 23.45
   + HMG 26.10 0.120 0.9440 0.9312 24.67
   + SMA 25.10 0.105 0.9380 0.9227 25.21
MILD (Full) 24.20 0.093 0.9703 0.9499 26.24
Table 2: Ablation study on the MILD dataset.
4.4 Compositional Flexibility of Layered Outputs
Beyond achieving high-quality foreground removal and background restoration, MILD’s layered diffusion design introduces a key advantage: the ability to flexibly recompose scenes using instance-specific foreground layers and a clean background. As shown in Figure 7, MILD generates disentangled outputs for each target, enabling selective combination to construct customized scenes. This allows users to separate specific individuals (e.g., only IP11), remove arbitrary targets (e.g., w/o IP33), or reconstruct any subset of the original scene. Such instance-aware recomposition offers fine-grained control over content manipulation, making MILD well-suited for interactive editing and scene simulation tasks. More details are shown in Appendix D.
Figure 7: Illustration of MILD’s compositional flexibility.
5 Conclusion
We present Multi-Instance Layered Diffusion, a diffusion-based strategy reformulating inpainting as a layered, decoupled generation process for precise human erasing, supported by our curated, high-quality MILD dataset. MILD leverages HMG to provide instance-aware structure understanding, and employs SMA to bring spatially-conditioned attention control, effectively reducing artifacts and improving reconstruction fidelity. Extensive experiments show that MILD outperforms existing inpainting baselines, achieving better semantic consistency, fewer artifacts, and more coherent scene reconstructions. MILD provides a promising direction for scaling precise diffusion-based inpainting to complex multi-object and open-domain removal tasks.
Appendix
Appendix A Dataset and Code Availability
Dataset Construction.
Our dataset comprises high-quality images collected from two primary sources: (1) a carefully curated subset of the OpenImages V5 dataset Kuznetsova et al. (2020), and (2) publicly available web images, gathered under fair use for academic research purposes. All images are manually verified to ensure diversity in scenes, human poses, occlusions, and background complexity.
Each image is annotated with pixel-level, instance-specific human masks. We further provide human-verified ground-truth completions to support supervised evaluation of object removal performance. Representative samples are shown in Figure 8. The complete dataset will be publicly released upon acceptance at: https://mild-multi-layer-diffusion.github.io/dataset/
Figure 8: Sample data collected from MILD dataset.
Code and Model Release.
To support future research and ensure reproducibility, we have created a public project page featuring representative examples and the paper abstract. The full codebase and pre-trained models will be made available upon acceptance at: https://mild-multi-layer-diffusion.github.io/
Appendix B More Related Works
Layered Generation and Controllability.
Recent works on image generation have explored layered modeling and layout-based generation to enhance object awareness and controllability. Layered modeling works Jia et al. (2024); Lu et al. (2020) leverage multi-level structures through latent decomposition and scene-layer separation to enable semantically consistent editing. However, they often operate within a unified generation pathway and lack explicit instance disentanglement or spatial conditioning, making them less suitable for multi-object removal. Prompt- and layout-controllable generation works Liu et al. (2022); Zheng et al. (2023); Zhangli et al. (2024) aim to synthesize scenes from object-level prompts or layout priors. Although these methods offer flexible conditioning and structural guidance, they generally operate at a holistic level without explicit separation of foreground instances, limiting their capacity to handle occlusions or support fine-grained instance control. In contrast, our approach introduces a mask-aware, multi-branch generation framework that explicitly disentangles each foreground instance and the background into independent pathways. This layered formulation enables precise object removal, instance-level recomposition, and enhanced semantic fidelity through joint supervision.
Appendix C More Framework Details
To ensure reproducibility, we provide additional details of the training pipeline, architectural design, and spatial supervision strategies used in LayeredUNet. The model is optimized in stages (Algorithm 2), where a training controller gradually activates different generation branches. Each image is encoded into the latent space and perturbed with spatial offsets derived from the union of all masks. To stabilize learning, the background branch is initially protected by freezing its gradients, while foreground branches are progressively optimized with instance-specific conditioning.
Both foreground and background branches share a unified UNet backbone but use independent LoRA adapters. During the forward pass (Algorithm 3), we apply Spatially-Modulated Attention (SMA) in all attention layers across both branches, replacing standard self-attention. SMA enforces spatial constraints guided by the input masks, helping to prevent feature leakage and preserve structural consistency.
To further improve boundary quality and compositional integrity, optional modules such as layer exchange and boundary smoothing can be enabled. These help align overlapping regions and reduce visible seams. Additionally, the foreground branch is supervised using a region-aware loss that emphasizes both the core masked region and its boundaries via residual and gradient-based penalties, improving local coherence and mask adherence.
Algorithm 2 Training of LayeredUNet with Progressive Dual LoRA
0: Input image $x$, mask set $\{M_{k}\}_{k=1}^{N}$, prompt $c_{\text{text}}$, optional pose $P$, parsing $S$
[…]
6:for each training step $t$ do
7:  $\text{stage}\leftarrow\text{GetTrainingStage}(t)$
8:  $z_{t}\leftarrow\text{AddNoise}(z_{0},t)$
9:  for $k=1$ to NN do
10:   $M_{\text{others}}^{(k)}\leftarrow\text{Aggregate}(\{M_{j}\}_{j\neq k})$
11:   $Z_{\text{mask}}^{(k)}\leftarrow\text{MaskEncoder}(M_{\text{others}}^{(k)})$
12:   $C_{k}\leftarrow\text{Concat}(Z_{\text{cond}},Z_{\text{mask}}^{(k)})$
13:   $\epsilon_{k}\leftarrow\text{UNet}(z_{t},t,C_{k},\text{branch}=\texttt{"fg"},M_{k})$
14:   $\mathcal{L}_{k}\leftarrow\text{RegionLoss}(\epsilon_{k},z_{0},M_{k})$
15:  end for
16:  $\epsilon_{\text{bg}}\leftarrow\text{UNet}(z_{t},t,Z_{\text{cond}},\text{branch}=\texttt{"bg"})$
17:  $\mathcal{L}_{\text{bg}}\leftarrow\|\epsilon-\epsilon_{\text{bg}}\|_{2}^{2}$
18:  $\mathcal{L}_{\text{total}}\leftarrow\lambda_{t}\sum_{k}\mathcal{L}_{k}+\mathcal{L}_{\text{bg}}$
19:  if stage requires gradient protection then
20:   FreezeLoRA("background")
21:  end if
22:  Backward($\mathcal{L}_{\text{total}}$), UpdateLoRA()
23:end for
Algorithm 3 Forward Pass of LayeredUNet with SMA
0: Latent input $z_{t}$, timestep $t$, condition $C$, mask $M$, branch $\in\{\texttt{"foreground"},\texttt{"background"}\}$
[…]
2:for each attention layer in UNet do
3:  if branch = "foreground" then
4:   $Q,K,V\leftarrow\text{LoRA}_{\text{fg}}(z_{t},C)$
5:   $\text{attn}\leftarrow\text{SMA}(Q,K,V)$
6:  else
7:   $Q,K,V\leftarrow\text{LoRA}_{\text{bg}}(z_{t},C)$
8:   $\text{attn}\leftarrow\text{SMA}(Q,K,V)$
9:  end if
10:  $z_{t}\leftarrow\text{Apply}(z_{t},\text{attn})$
11:  if layer exchange enabled then
12:   $z_{t}\leftarrow z_{t}+\text{LayerExchange}(z_{t})$
13:  end if
14:  $z_{t}\leftarrow z_{t}+\text{BoundarySmoothing}(z_{t},M)$
15:end for
16:return  $\epsilon_{\text{pred}}\leftarrow\text{UNetOutput}(z_{t})$
Appendix D More Experiments
D.1 More Quantitative Results
To further validate the robustness and generalizability of our method, we introduce the Structural Similarity Index Measure (SSIM) as an additional evaluation metric. Unlike FID, LPIPS, or CLIP-based perceptual scores, SSIM captures low-level structural consistency between the generated and ground-truth images, providing complementary insight into image fidelity.
As summarized in Table 3 (MILD dataset) and Table 4 (OpenImages dataset), our method consistently outperforms prior approaches across all major metrics, including the newly introduced SSIM. This further confirms that MILD not only produces perceptually realistic outputs but also preserves structural coherence with the surrounding context.
The inclusion of SSIM highlights MILD’s ability to maintain both global semantics and local structural details, reinforcing its effectiveness in complex object removal scenarios.
[…]
SDXL Inpaint 53.45 0.242 0.8398 0.8331 18.22 0.6984
PowerPaint 48.04 0.207 0.8901 0.8668 20.27 0.7569
LaMa 35.02 0.164 0.8919 0.8872 24.13 0.7984
CLIPAway 48.86 0.230 0.8647 0.8605 19.49 0.7046
RoRem 40.41 0.196 0.8888 0.8810 22.35 0.7521
MILD (Ours) 24.20 0.093 0.9703 0.9499 26.24 0.8341
Table 3: Quantitative comparison of MILD and other methods on MILD dataset. Best results are bold, second best are underlined.
[…]
SDXL Inpaint 26.22 0.1346 0.8741 0.8751 21.88 0.8290
PowerPaint 10.56 0.0488 0.9649 0.9678 31.04 0.9638
Inst-Inpaint 11.42 0.4100 0.7400 0.8207 23.75 0.8023
LaMa 10.38 0.0470 0.9192 0.9293 31.83 0.9461
CLIPAway 22.73 0.1283 0.8973 0.8985 23.59 0.8276
RoRem 18.23 0.0982 0.9095 0.9148 26.59 0.8713
MILD (Ours) 17.86 0.0668 0.9829 0.9700 32.14 0.9515
Table 4: Quantitative comparison of object removal methods on the OpenImages dataset. Best results are bold, second best are underlined.
D.2 More Qualitative Results
To further demonstrate the effectiveness of our approach, we present additional qualitative comparisons against state-of-the-art baselines in diverse and challenging scenarios. These include scenes with multiple interacting persons, complex occlusions, cluttered backgrounds, and fine structural details such as limbs, shadows, and object boundaries.
As shown in Figure 9, our method produces more semantically consistent and visually coherent results compared to existing methods. In particular, our model excels at generating plausible textures, maintaining global scene structure, and avoiding common failure cases such as blurry artifacts, object duplication, or semantic distortions. These results further validate the benefits of our layered generation strategy, human-centric guidance, and attention modulation.
All examples are selected from the test split of our curated dataset. Please refer to the project page for high-resolution images and interactive comparisons.
Figure 9: More Qualitative Comparisons between MILD and other methods on MILD dataset.
D.3 Inference Efficiency.
Table 5 summarizes the average inference time per image at $512\times 512$ resolution using an A800 GPU with 25 diffusion steps. Despite incorporating a layered, multi-branch generation strategy, MILD maintains efficient runtime performance. Due to the shared UNet backbone and lightweight LoRA-based adaptation, our method achieves instance-aware generation without redundant computation overhead.
Compared to other diffusion-based baselines such as PowerPaint, CLIPAway, and RoRem, MILD demonstrates notably faster inference. While it supports text prompts to enhance generation precision, the impact on speed is minimal, resulting in comparable runtime to the SDXL Inpainting baseline. Although LaMa achieves the fastest runtime overall owing to its CNN-based architecture, it compromises significantly on visual fidelity and semantic accuracy, as evidenced by its performance in Table 4. Overall, MILD offers a favorable trade-off between generation quality and runtime efficiency.
Table 5: Average inference time per image (in seconds), along with framework type and guidance type.
Method Guidance Type Time (s)
SDXL Inpaint Text + Mask 1.79
PowerPaint Text + Mask 6.21
Inst-Inpaint Text-only 2.21
LaMa Mask-only 1.50
CLIPAway Text + Mask 4.76
RoRem Text + Mask 4.15
MILD (Ours) Text + Mask 1.91
D.4 AI evaluation.
Implementation details
We evaluate perceptual quality using both AI-based and human-based protocols. For the AI evaluation, we employ GPT-4o to assess 100 randomly selected image pairs. Given a standardized prompt, GPT-4o performs a side-by-side visual inspection of the original and edited images and assigns four aspect-wise scores including naturalness, semantic consistency, visual harmony, and artifact suppression, each on a 1–10 scale. These are then averaged to produce an overall perceptual score. The model also provides a binary decision indicating whether the object removal is successful or unsuccessful.
For the human evaluation, we recruit 20 participants to rate the same 100 image pairs. Each participant is asked to consider the above four aspects holistically and assign a single overall quality score between 1 and 10. Additionally, they indicate whether the result is successful or unsuccessful, based on the overall plausibility and absence of noticeable artifacts. Final human scores are computed by averaging all participant ratings, and success rate is calculated as the percentage of participants marking each result as successful.
AI & Human Evaluation Protocol
To assess the perceptual quality of human erasing, we employ a unified evaluation protocol involving both AI-based and human-based assessments. In both cases, evaluators are presented with side-by-side comparisons of the original image (with the human present) and the edited result, and are asked to judge the visual plausibility of the removal without access to ground truth.
For the AI evaluation, we use GPT-4o to score each image across four criteria, adapted from standard visual editing assessments: naturalness, semantic consistency, visual harmony, and artifact suppression. Each aspect is rated on a 1–10 scale, and the average of these scores is reported as the overall perceptual score. Additionally, GPT-4o provides a binary decision indicating whether the removal is considered successful or unsuccessful, based on visual realism and contextual coherence.
  • Naturalness: Whether the filled region appears visually plausible and free from obvious artifacts.
  • Semantic Consistency: Whether the inpainted content aligns semantically with the surrounding scene.
  • Visual Harmony: Whether lighting, textures, and shadows in the edited region are consistent with the rest of the image.
  • Artifact Suppression: Whether the result avoids seams, blurs, distortions, or other visual artifacts.
In the human evaluation, we recruit 20 participants to assess the same set of 100 images. Unlike the AI annotator, human raters do not score each aspect individually. Instead, they are instructed to consider all four criteria jointly and provide a single overall quality score (1–10) for each image, reflecting their holistic impression. They also indicate whether the removal is successful or unsuccessful according to the same definition used in the AI evaluation—namely, that a successful result should appear seamless, contextually appropriate, and free of noticeable flaws.
Final results are reported in terms of AI and human perceptual scores, along with success rates computed as the percentage of cases marked as successful.
Evaluation Results
Table 6 summarizes the perceptual evaluation results from both AI and human annotators. Our method, MILD, consistently achieves the highest overall perceptual score among all compared approaches. It also attains the highest success rates from GPT-4o and from human raters, indicating strong agreement across evaluation modalities. Overall, these results confirm the effectiveness of our MILD strategy in producing perceptually convincing and contextually coherent removal outputs, outperforming prior methods in both automated and human assessments.
Method AI Overall Human Overall Success Rate
SDXL Inpaint 3.7 4.2 53% / 50%
PowerPaint 5.0 5.2 58% / 54%
Inst-Inpaint 4.7 4.8 48% / 41%
LaMa 6.4 6.3 67% / 63%
CLIPAway 3.9 4.0 52% / 49%
RoRem 5.4 5.4 64% / 61%
MILD (Ours) 8.3 8.4 82% / 78%
Table 6: Perceptual evaluation summary on the MILD dataset. AI and human scores reflect overall perceptual quality (1–10 scale).
Figure 10: Compositional scene recomposition with MILD. Given a masked input with three foreground targets (IP11–IP33), we show the disentangled outputs and six representative recombinations enabled by our layered generation design.
D.5 More Compositional Flexibility
Figure 10 provides a detailed visualization of MILD’s compositional flexibility. Given a masked input image containing three interacting persons (IP11–IP33), our model generates four disentangled outputs: one foreground layer per instance and a clean background. By selectively combining these layers, MILD enables the construction of diverse scene variants. As illustrated, we showcase six recomposed outputs, including single-subject isolation (e.g., only IP11), targeted removal (e.g., w/o IP22), and partial restoration (e.g., IP11 and IP33). This layered formulation supports fine-grained, user-controllable editing, and highlights the advantages of explicit instance-level generation over unified inpainting approaches.
Appendix E Limitations
Despite the strong performance of MILD in a wide range of challenging scenarios, several limitations remain. First, when the occluded region is particularly large, occupying up to two-thirds of the image, MILD may struggle to produce fully realistic completions. While the results are generally more perceptually plausible than those generated by existing baselines (see Figure 11, Row 1), the reconstruction of fine-grained background details is often imperfect due to limited contextual information. Second, in some instances, the generation produced by MILD tends to be overly conservative. Although the object is successfully removed and the result appears semantically coherent, the model may fail to fully restore the background content or reflect the underlying scene logic (see Figure 11, Row 2).
We consider these issues valuable directions for future research, particularly in enhancing hallucination capabilities under severe occlusion and developing more flexible generation mechanisms that better balance fidelity and expressiveness.
Figure 11: Failure example of MILD, along with other methods.